# Neural Attention Search

### v2

@@ 5 Conclusion and Future Work
Efficiently managing the KV cache is crucial for deploying large language models in long-context applications. In this work, we propose NAtS, an approach that automatically optimizes the optimal roles of each token to determine how long they should survive. By constructing a learnable mask, NAtS learns to sparsify the attention map end-to-end. Our experiments show that NAtS could use much less KV caches compared to the State-of-the-art KV caching reduction approach. While NAtS shows promising results, future work could include exploration of further token roles and structured search spaces with hierarchies. Overall, we believe that NAtS paves the way towards more efficient and scalable inference with LLMs.
6 Impact Statements
@@ 7 Acknowledgement
Appendix A Details on Backward Propagation
A.1 Gradients for Attention Masks
@@
In Eq. 15 and  16, we show the gradient for $M$ is the same as the value that $\partial O/\partial A$ is supposed to be if no mask is applied. Since we have $g_{i,j}=e^{A_{i,j}}\odot M_{i,j}$. Let’s set $S_{i}:=\sum_{j}g_{i,j}$ and $P_{i,j}:=\frac{g_{i,j}}{S_{i}}$. Then we have $\mathbf{d}\mathbf{P}=\mathbf{d}\mathbf{O}\mathbf{V}^{T}$. Therefore,
@@
Following FlashAttention (Dao et al., 2022), we define $D_{i}=do_{i}^{T}o_{i}$, then
$dM=\frac{e^{A_{i,j}}}{S_{i:}}(dP_{ij}-D_{i})$ (20)
$dA_{i,j}=P_{i,j}(dP_{ij}-D_{i})$ (21)
@@
However, in practice, we found this makes the training unstable. All the values in a row with vanilla softmax attention are normalized such that their sum is bounded. However, we need to use the unmasked $e^{M}$ to compute $dM$ in our scenerio, meaning $\frac{e^{A_{i,j}}}{S_{i}}$ could be unbounded. This might result in an unstable training process. To alleviate this, we record both the masked denominator $S_{i}$ and unmasked denominator $S^{\prime}_{i}=\sum_{j}e^{A_{i,j}}$ during the forward pass. We then use the unmasked denominator to compute the unmasked $P^{\prime}_{i,j}=\frac{e^{A_{i,j}}}{S^{\prime}_{i}}$ and further apply $p^{\prime}$ to compute $dM$:
$dM=P^{\prime}_{i,j}(dP_{ij}-D_{i})$ (22)
This only requires storing an additional $S^{\prime}_{i}$ during the forward pass (with a complexity of $O(\mathit{L})$) and one additional computation during the backward pass (since $S^{\prime}_{i}$ is already computed and stored during the forward pass).
A.2 Details on computing $d\alpha^{G-}_{i}$
In Eq. 11 and  12, we show that an additional $d\alpha^{G-}_{i}$ needs to be computed for each token’s gradient for Global Tokens. Figure  14 illustrates an example of this. Token 4 is a global token, and we search for its next Global Token (which is Token 10 in this example). Assuming that we want to change Token 4 to another role, the regions within the boundary (orange ones) are those tokens that are influenced by this swtching: given that Token 4 no longer becomes Global Token, Token 2 and 3 will be part of a larger subsequence that ranges from 1 to 10, and the attentino maps within the orange region should not be masked out. Intuitively, this gradient measures the regret that we made in order to switch one Global Token into a Local Token. A similar idea can be found in the red region. Assuming that we want to switch Token 7 to a Global Token, then Tokens 5,6 will be split into a new subsequence and we will no longer connect them with Token 8, 9, 10 since they belong to two different sub-sequences after the switch. Hence, this value in the red regions measures the regret if we mistakenly classify a Global Token as non-Global Tokens.
@@ A.2 Details on computing $d\alpha^{G-}_{i}$
In practice, the $d\alpha^{G-}_{i}$ values for Global Token can be easily collected since we can easily get the last and next $G$ index with a scan function and collect the gradients to the corresponding positions. However, for those values on non Global Tokens, since the entire sequence can be a sub-sequence if no Global Token exists in the current sequence, collecting all the values that are below and on the left side of the Local Tokens might be too expensive. Hence, we ask each Local Token can only visit a fixed number of tokens on both sides. In this case, this operation is equivalent to a convolutional operation with weights of either 1 or 0. We set this value as 16 in our experiments to fit the grid sizes in triton.
A.3 Optimizing for sparser Attention Maps
The sparse regularization term $\lambda$ is directly applied to the corresponding gradients for the Global Tokens and Local Tokens. This value should penalize the number of unmasked tokens for each row. While the number of unmasked tokens for Global Token and Local Tokens in column $i$ are $\mathit{L}-i$ and $K_{i}-i$ with $K_{i}$ defined in Section 3.2. Hence, we have:
@@ A.3 Optimizing for sparser Attention Maps
$dG^{L_{sparse}}_{i}$ $=\lambda\times\frac{K_{i}-i}{\mathit{L}}$ (24)
Combining Eq.  9,  11,  23 and Eq.  24, we have:
@@
with $M_{i,j}^{G},M_{i,j}^{SW},M_{i,j}^{L}$ defined in Eq.  4,  7 and  5 and $M_{i,j}^{casual}$ is a casual attention mask.
Appendix B Experiments Details
B.1 Collecting the Fine-tune Training Set

### v3

@@ 5 Conclusion and Future Work
Efficiently managing the KV cache is crucial for deploying large language models in long-context applications. In this work, we propose NAtS, an approach that automatically optimizes the optimal roles of each token to determine how long they should survive. By constructing a learnable mask, NAtS learns to sparsify the attention map end-to-end. Our experiments show that NAtS uses much less KV caches compared to the State-of-the-art KV caching reduction approach. While NAtS shows promising results, future work could include exploration of further token roles and structured search spaces with hierarchies. Overall, we believe that NAtS paves the way towards more efficient and scalable inference with LLMs.
@@ 6 Acknowledgement
Appendix A Limitations
While NAtS shows promising results that efficiently reduce the KV cache sizes and our search space already contains most of the KV cache eviction strategies. However, the construction of the attention mask is still column-wise oriented and mostly focuses on the vertical style sparse attention [41]. Hence, this work does not involve the block-sparse attention and slash attention that are widely observed in other frameworks [41, 48]. One potential future direction would be to introduce a more structured search space that allows for the construction of more sparse attention types.
Appendix B Broader Impact
Appendix C Details on Backward Propagation
C.1 Gradients for Attention Masks
@@
In Eq. 9 and  10, we show the gradient for $M$ is the same as the value that $\partial O/\partial A$ is supposed to be if no mask is applied. Since we have $g_{i,j}=e^{A_{i,j}}\odot M_{i,j}$. Let’s set $S_{i}:=\sum_{j}g_{i,j}$ and $P_{i,j}:=\frac{g_{i,j}}{S_{i}}$. Then we have $\mathbf{d}\mathbf{P}=\mathbf{d}\mathbf{O}\mathbf{V}^{T}$. Therefore,
@@
Following FlashAttention [18], we define $D_{i}=do_{i}^{T}o_{i}$, then
$dM_{i,j}=\frac{e^{A_{i,j}}}{S_{i:}}(dP_{ij}-D_{i})$~ (14)
$dA_{i,j}=P_{i,j}(dP_{ij}-D_{i})$~ (15)
@@
However, in practice, we cannot guarantee an upper bound for $\frac{e^{A_{i,j}}}{S_{i}}$ with $M_{i,j}=0$ since $S_{i}:=\sum_{j}e^{A_{i,j}}\odot M_{i,j}$. As a result, $\lim_{e^{(A_{i,j})}\to\infty}(\frac{e^{(A_{i,j})}}{S_{i}})=\infty$ when $M_{i,j}=0$. Hence, we first clip $\frac{e^{A_{i,j}}}{S_{i}}$ within $(0,1)$ and then compute $dM$ with the clipped value:
$P^{\prime}_{i,j}$ $=\min(\frac{e^{A_{i,j}}}{S_{i}},1)$ (16)
$dM_{i,j}$ $=P^{\prime}_{i,j}(dP_{ij}-D_{i})$~ (17)
C.2 Details on computing the gradients for computing the token gradients info $d\alpha$
$d\alpha_{i}=\sum_{j}dM_{i,j}\times M_{i,j}^{\alpha}$~ (18)
Where $\alpha\in\{G,L,SW\}$ is the discretized token type. Intuitively, this shows the model’s preference over short-range or long-range correlations: If $i$ is quite close to $j$, then all the $\alpha$ will receive the same gradient information. However, if the model wants to create a long-range correlation with $i\gg j$, only Global Tokens will receive the gradient information. The network will therefore prefer to classify the corresponding tokens as global tokens.
In cases where $G_{k}$ is 0, Equation 19 is the negative value of Equation 5. However, for the case where $G_{k}=1$, this is equivalent to the local mask values where $k$ is no longer set as a Global Token. This requires us to find the index of the next global token $I^{G_{\text{next}}}_{k+1}$ and the last global token $I^{G_{\text{next}}}_{k-l}$ where $l$ is the length of the local sequence that ends at $k$.
$d\alpha^{G-}_{i}$ $=\begin{cases}\sum\limits_{\begin{subarray}{c}m\geq i\\ i>n\geq I^{G_{\text{next}}}_{k-l}\end{subarray}}M^{L}_{m,n}\times dM_{m,n}\ \ &\text{if}\ G_{i}=0,\\ \sum\limits_{\begin{subarray}{c}I^{G_{\text{next}}}_{i+1}\geq m>=i\\ i>n\geq I^{G_{\text{next}}}_{k-l}\end{subarray}}dM_{m,n}\ \ &\text{if}\ G_{i}=1,\end{cases}$ (20)
$d\alpha^{G}_{i}$ $=d\alpha^{G}_{i}-d\alpha^{G-}_{i}$ (21)
Figure  8 illustrates an example of this. Token 4 is a global token, and we search for its next Global Token (which is Token 10 in this example). Assuming that we want to change Token 4 to another role, the regions within the boundary (orange ones) are those tokens that are influenced by this swtching: given that Token 4 no longer becomes Global Token, Token 2 and 3 will be part of a larger subsequence that ranges from 1 to 10, and the attentino maps within the orange region should not be masked out. Intuitively, this gradient measures the regret that we made in order to switch one Global Token into a Local Token. A similar idea can be found in the red region. Assuming that we want to switch Token 7 to a Global Token, then Tokens 5,6 will be split into a new subsequence and we will no longer connect them with Tokens 8, 9, 10 since they belong to two different sub-sequences after the switch. Hence, this value in the red regions measures the regret if we mistakenly classify a Global Token as non-Global Tokens.
@@ C.2 Details on computing the gradients for computing the token gradients info $d\alpha$
In practice, the $d\alpha^{G-}_{i}$ values for Global Token can be easily collected by looking at the last and next $G$ index with a scan function to collect the gradients to the corresponding positions. However, computing $d\alpha^{G-}_{i}$ with $G_{i}=0$ requires the gradient information from all the past token mask information. This might be computationally prohibitive in practice, since the valid values contained in Equation 20 are sparse, we ignore the cases for $G_{i}=0$ and only compete the dense gradeints for the case where $G_{i}=1$.
C.3 Optimizing for sparser Attention Maps
The sparse regularization term $\lambda$ is directly applied to the corresponding gradients for the Global Tokens and Local Tokens. This value should penalize the number of unmasked tokens for each row. While the number of unmasked tokens for Global Token and Local Tokens in column $i$ are $\mathit{L}-i$ and $I^{G_{\text{next}}}_{i}-i$ with $I^{G_{\text{next}}}_{i}$ defined in Section 3.2. Hence, we have:
@@ C.3 Optimizing for sparser Attention Maps
$dG^{L_{sparse}}_{i}$ $=\lambda\times\frac{I^{G_{\text{next}}}_{i}-i}{\mathit{L}}$ (23)
Combining Eq.  18,  20,  22 and Eq.  23, we have:
@@
with $M_{i,j}^{G},M_{i,j}^{SW},M_{i,j}^{L}$ defined in Eq.  3,  6 and  4 and $M_{i,j}^{casual}$ is a casual attention mask.
Appendix D Integrating NAtS into FlashAttention
Algorithm 1 NAtS forward pass on FlashAttention2, we mark the  NAtS related operations with blue
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{N\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W
1: Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$, $\mathbf{\alpha},\mathbf{I^{G_{\text{next}}}}$ into $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$, $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$, $\mathbf{\alpha}_{1},\dots,\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ , and $\mathbf{I^{G_{\text{next}}}}_{1},\dots,\mathbf{I^{G_{\text{next}}}_{T_{c}}}$ of size $B_{r}$ each.
2: Divide the output $\mathbf{O}\in\mathbb{R}^{N\times d}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, and divide the logsumexp $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3:for $1\leq i\leq T_{r}$ do
4:   Load $\mathbf{Q}_{i}$ from HBM to on-chip SRAM.
5:   On chip, initialize $\mathbf{O}_{i}^{(0)}=(0)_{B_{r}\times d}\in\mathbb{R}^{B_{r}\times d},\ell_{i}^{(0)}=(0)_{B_{r}}\in\mathbb{R}^{B_{r}},m_{i}^{(0)}=(-\infty)_{B_{r}}\in\mathbb{R}^{B_{r}}$.
6:  for $1\leq j\leq T_{c}$ do
7:   Load $\mathbf{\alpha}_{j}$ and $\mathbf{I^{G_{\text{next}}}}_{j}$ to SRAM and check if the current block contain any valid values according to Equations 3,4,6,
8:   if Do Compute then
9:    Construct attention mask $\mathbf{M}_{i,j}$ with $\mathbf{\alpha}_{j}$, $\mathbf{I^{G_{\text{next}}}}_{j}$, and W with Equations 3,4,6
10:     Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to on-chip SRAM.
11:     On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
12:     On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$, $\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\odot\mathbf{M}_{i,j}}\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
13:     On chip, compute $\mathbf{O}_{i}^{(j)}=\mathrm{diag}(e^{m_{i}^{(j-1)}-m_{i}^{(j)}})^{-1}\mathbf{O}_{i}^{(j-1)}+\tilde{\mathbf{P}}_{i}^{(j)}\mathbf{V}_{j}$.
14:   end if
15:  end for
16:  On chip, compute $\mathbf{O}_{i}=\mathrm{diag}(\ell_{i}^{(T_{c})})^{-1}\mathbf{O}_{i}^{(T_{c})}$.
17:  On chip, compute $L_{i}=m_{i}^{(T_{c})}+\log(\ell_{i}^{(T_{c})})$.
18:  Write $\mathbf{O}_{i}$ to HBM as the $i$-th block of $\mathbf{O}$.
19:  Write $L_{i}$ to HBM as the $i$-th block of $L$.
20:end for
21: Return the output $\mathbf{O}$ and the logsumexp $L$.
Flashattention [17] continuously loads Q, K, V values to the SRAM and only computes the attention maps within the loaded blocks.  NAtS loads the corresponding token states information $\alpha$ and construct the masks $M$ dynamically with Equation 3,  5,  4. During the backward process, Flashattention collects the row-wise gradients for computing $d\mathbf{Q}$ and column-wise gradients to compute $d\mathbf{K}$ and $d\mathbf{V}$. Hence, the gradients $d\mathbf{\alpha}$ can be computed together with $d\mathbf{K}$ and $d\mathbf{V}$. We show how  NAtS can be integrated into Flashattion2  [17] in Algorithms 1 and 2. We highlight the difference between the vanilla Flashattention2 and Flashattention2 with  NAtS in blue. Compared with vanilla Flashattention models,  NAtS adaptively skips the blocks that do not have valid mask values to avoid unnecessary computation. These invalid blocks will only be applied to update the gradients for $d\mathbf{\alpha}$.
Algorithm 2 NAtS backward pass on FlashAttention2, we mark the  NAtS related operations with blue
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{O},\mathbf{dO}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{\mathit{L}\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{\mathit{L}}$ in HBM, vector $L\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W.
1: Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$, $\mathbf{\alpha},\mathbf{I^{G_{\text{next}}}}$ in to $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$ and $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$, $\mathbf{\alpha}_{1},\dots,\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ , and $\mathbf{I^{G_{\text{next}}}}_{1},\dots,\mathbf{I^{G_{\text{next}}}_{T_{c}}}$ of size $B_{r}$ each.
2: Divide $\mathbf{O}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, divide $\mathbf{dO}$ into $T_{r}$ blocks $\mathbf{dO}_{i},\dots,\mathbf{dO}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3: Initialize $\mathbf{dQ}=(0)_{N\times d}$ in HBM and divide it into $T_{r}$ blocks $\mathbf{dQ}_{1},\dots,\mathbf{dQ}_{T_{r}}$ of size $B_{r}\times d$ each. Divide $\mathbf{dK},\mathbf{dV}\in\mathbb{R}^{N\times d}$ in to $T_{c}$ blocks $\mathbf{dK}_{1},\dots,\mathbf{dK}_{T_{c}}$ and $\mathbf{dV}_{1},\dots,\mathbf{dV}_{T_{c}}$, of size $B_{c}\times d$ each. Divide $d\mathbf{\alpha}\in\mathbb{R}^{\mathit{N_{\text{opts}}}}$ in to $d\mathbf{\alpha}_{1},\dots,d\mathbf{\alpha}_{T_{c}}$ of size $B_{r}\times\mathit{N_{\text{opts}}}$ each.
4: Compute $D=\mathrm{rowsum}(\mathbf{dO}\odot\mathbf{O})\in\mathbb{R}^{d}$ (pointwise multiply), write $D$ to HBM and divide it into $T_{r}$ blocks $D_{1},\dots,D_{T_{r}}$ of size $B_{r}$ each.
5:for $1\leq j\leq T_{c}$ do
6:  Load $\mathbf{K}_{j},\mathbf{V}_{j}$, $\mathbf{\alpha}_{j}$ and $\mathbf{I^{G_{\text{next}}}}_{j}$ from HBM to on-chip SRAM.
7:  Initialize $\mathbf{dK}_{j}=(0)_{B_{c}\times d},\mathbf{dV}_{j}=(0)_{B_{c}\times d}$, $d\mathbf{\alpha}_{j}=(0)_{B_{c}\times\mathit{N_{\text{opts}}}}$ on SRAM.
8:  for $1\leq i\leq T_{r}$ do
9:   Load $\mathbf{Q}_{i},\mathbf{O}_{i},\mathbf{dO}_{i},\mathbf{dQ}_{i},L_{i},D_{i}$ from HBM to on-chip SRAM.
10:   On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
11:   On chip, Construct attention mask $\mathbf{M}_{i,j}$ with $\mathbf{\alpha}_{j}$, $\mathbf{I^{G_{\text{next}}}}_{j}$, and W with Equations 3,4,6
12:   On chip, compute $\mathbf{P^{\prime}}_{i}^{(j)}=\exp(\mathbf{S}_{ij}-L_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$ and $\mathbf{P}_{i}^{(j)}=\mathbf{P^{\prime}}_{i}^{(j)}\odot\mathbf{M}_{i,j}$.
13:   Check if the current block contain any valid value
14:   if Do compute then
15:    On chip, compute $\mathbf{dV}_{j}\leftarrow\mathbf{dV}_{j}+(\mathbf{P}_{i}^{(j)})^{\top}\mathbf{dO}_{i}\in\mathbb{R}^{B_{c}\times d}$.
16:    On chip, compute $\mathbf{dP}_{i}^{(j)}=\mathbf{dO}_{i}\mathbf{V}_{j}^{\top}\in\mathbb{R}^{B_{r}\times B_{c}}$.
17:    On chip, compute $\mathbf{dS}_{i}^{(j)}=\mathbf{P}_{i}^{(j)}\odot(\mathbf{dP}_{i}^{(j)}-D_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$.
18:    Load $\mathbf{dQ}_{i}$ from HBM to SRAM, then on chip, update $\mathbf{dQ}_{i}\leftarrow\mathbf{dQ}_{i}+\mathbf{dS}_{i}^{(j)}\mathbf{K}_{j}\in\mathbb{R}^{B_{r}\times d}$, and write back to HBM.
19:    On chip, compute $\mathbf{dK}_{j}\leftarrow\mathbf{dK}_{j}+{\mathbf{dS}_{i}^{(j)}}^{\top}\mathbf{Q}_{i}\in\mathbb{R}^{B_{c}\times d}$.
20:   end if
21:   On chip, clip $\mathbf{P^{\prime}}_{i}^{(j)}$ to (0,1) and compute $d\mathbf{S^{\prime}}_{i}^{(j)}=\mathbf{P^{\prime}}_{i}^{(j)}\odot(\mathbf{dP}_{i}^{(j)}-D_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$
22:   On chip, update $d\mathbf{\alpha}_{j}$ with Equations 24, 25, 26
23:  end for
24:  Write $\mathbf{dK}_{j},\mathbf{dV}_{j}$, and $d\mathbf{\alpha}_{j}$ to HBM.
25:end for
26: Return $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$.
Appendix E Experiments Details
Here we discuss further details in our experiments. The codes for  NAtS can be found under  https://github.com/automl/NeuralAttentionSearch
E.1 Detailed Hyperparameter Setting for GPT2-small Training
Following the setting from NanoGPT [43], the GPT-2 style small has 12 layers and 12 heads with a hidden dimension of 768. Instead of the learnable position encoding, we apply rotary embeddings [72] to each transformer layer. The PG-19 dataset contains books extracted from Project Gutenberg [65] with about 2B tokens in the training sets. We train all models with a context length of 1024 and a batch size of 480 (using gradient accumulation). We train them for 600 000600\,000 iterations evaluate them on the test sets of PG19 with a context length of 1024. Training one model can be finished in 16 hours with 4 Nvidia H100 PCEi GPUs.
E.2 Collecting the Fine-tune Training Set